You invested $1,300,000 with a market-neutral hedge fund manager. The fee structure is 2/20, and the fund has a high-water-mark provision. Suppose the first year the fund manager loses 5 percent and the second year she gains 17 percent. Assume management fees are paid at the beginning of each year and performance fees are taken at the end of each year. What are the management and performance fees paid each year? (Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations. Round your answers to 2 decimal places.)

Year 1 Year 2

Management Fees $

Performance Fees $
Answer #1

Year 1:

Management Fee = Amount Invested * Management Fee% = $1,300,000 * 0.02 = $26,000

Performance Fee = $0; As there is no performance fee since the manager had a negative return.

Year 2:

Balance at the end of 1st year = Amount Invested - Management Fee for the 1st year

= $1,300,000 - $26,000 = $1,274,000

Management Fee = [Balance at the end of 1st year * (1 + Return in 1st year)]* Management Fee%

= [$1,274,000 * {1 + (-0.05)}] * 0.02 = $1,210,300 * 0.02 = $24,206

Balance at the end of 2nd year = [Balance at the end of 1st year - Management Fee for Year 1] * [1 + Return in 2nd year]

= [$1,274,000 - $24,206] * [1 + 0.17] = $1,249,794 * 1.17 = $1,462,258.98

Performance Fee = [Balance at the end of 2nd year - Watermark Level] * Performance Fee%

= [$1,462,258.98 - $1,300,000] * 0.20

= $162,258.98 * 0.20

= $32,451.80